Question:

A, B, C try to hit a target simultaneously but independently. Their respective probabilities of hitting the targets are \(\frac{3}{4},\frac{1}{2},\frac{5}{8}\). The probability that the target is hit by A or B but not by C is :

Show Hint

Events are independent, so multiply P(A or B) by P(C misses).
Updated On: Oct 31, 2026
  • 21/64
  • 7/8
  • 7/32
  • 9/64
Show Solution
collegedunia
Verified By Collegedunia

The Correct Option is A

Solution and Explanation

\(P(A\text{ or }B)=1-P(\bar A)P(\bar B)=1-\dfrac14\cdot\dfrac12=\dfrac78\).
\(P(\bar C)=1-\dfrac58=\dfrac38\).
By independence, required probability \(=\dfrac78\cdot\dfrac38=\dfrac{21}{64}\).
Answer: (1) 21/64
Was this answer helpful?
0
0